All of the resistors in this circuit have a resistance of R = 3.00Ω. What is the current, voltage drop, and power consumed in the resistor at the very bottom?

30 für 30 +10v +v 30 LW
Answer #1

here, for the net resistance in the circuit to the right

as 3 ohm and 3 ohm in parallel

R1 = 3/2 = 1.5 Ohm

for 1.5 and 3 ohm in series

Req = 3 + 1.5 = 4.5 Ohm

current in the 3 ohm resistance = 5/4.5 = 1.11 A

current in the bottom 3 Ohm = 1.11/2 = 0.56 A

voltage drop across the bottom 3 Ohm = 0.56 * 3

voltage drop across the bottom 3 Ohm = 1.67 V

the voltage drop across the bottom 3 Ohm is 1.67 V

Power consumed in the resistor = I^2 * R

Power consumed in the resistor = 0.56^2 * 3

Power consumed in the resistor = 0.941 W